OpenStudy (anonymous):

Hey! Describe the roots of the equation shown below: 3x^2-10x-5

OpenStudy (skullpatrol):

Hey! is for horses, saying "please" may get you help.

OpenStudy (skullpatrol):

Do you know the FOIL method of multiplying binomials?

OpenStudy (skullpatrol):

First of all that is an "expression" and it is not an "equation." Can you tell me why?

OpenStudy (anonymous):

But obviously it isn't.

OpenStudy (skullpatrol):

Correct. An equation MUST have an equals sign.

OpenStudy (skullpatrol):

So what is it suppose to equal???

OpenStudy (anonymous):

God I'm so dumb. Sorry. It's supposed to equal 0. I didn't see that I missed that while typing

OpenStudy (skullpatrol):

np, so $$3x^2-10x-5=0$$

OpenStudy (anonymous):

Correct!

OpenStudy (skullpatrol):

Ok, what are the factors of 3?

OpenStudy (anonymous):

I'm not sure... Would it be the other numbers in the equation?

OpenStudy (skullpatrol):

what two numbers multiplied together give 3?

OpenStudy (unklerhaukus):

For a quadratic equation\[ax^2+bx+c=0\] the roots can be found using the quadratic formula \[x_{1,2}=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

Still not sure. 3*1?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

Phew thought it was a trick question

OpenStudy (skullpatrol):

and how about 5?

OpenStudy (anonymous):

3*1.66?

OpenStudy (skullpatrol):

what two numbers multiply together to give 5?

OpenStudy (anonymous):

OH. 5*1.

OpenStudy (skullpatrol):

put those factors into $$3x^2-10x-5=(ax+b)(cx+d)$$ and usee FOIL until you get the right answer

OpenStudy (anonymous):

Could you put those factors into $$3x^2-10x-5=(ax+b)(cx+d)$$ so I can see where they are placed?

OpenStudy (ikram002p):

the solution is not integer u can describe the roots like this there exist c belong to (0,-1) { u might find onother interval } such that 3c^2-10c-5

OpenStudy (skullpatrol):

@mathhelpplease12 @ikram002p is right

OpenStudy (anonymous):

Oh okay. Thank you guys so much for the help!

OpenStudy (skullpatrol):

thanks for asking

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=3x^2-10x-5%3D0 This is to check your answer. You can practice factoring polynomial at http://www.saab.org/mathdrills/factor.cgi

OpenStudy (campbell_st):

wow... I thought this question would require the use of the discriminant... \[\Delta = b^2 - 4ac\] as the question says..."Describe the roots.... would have saved a lot of work... but hey... reading the question is the biggest part of answering a question